OpenStudy (anonymous):

square root of -7 and the square root of -7 equals what?

OpenStudy (anonymous):

49

OpenStudy (anonymous):

-7...

OpenStudy (anonymous):

its 49

OpenStudy (anonymous):

If this is the question \[\sqrt{-7}\sqrt{-7}\]

OpenStudy (anonymous):

time each other yes

OpenStudy (anonymous):

7

OpenStudy (anonymous):

\[\sqrt{7}(i ^{2})=-7\]

OpenStudy (anonymous):

oops that should be a square root of 7 squared

OpenStudy (anonymous):

oops that should be a square root of 7 squared\[(\sqrt{7})^{2}i ^{2}=-7\]

OpenStudy (anonymous):

oh ok thanks

OpenStudy (anonymous):

\[\sqrt{-7}*\sqrt{-7}\] \[\sqrt{(-7)(-7)}\] \[\sqrt{49}\] =7

OpenStudy (anonymous):

no corec...

OpenStudy (anonymous):

take i out of the radical first.

OpenStudy (anonymous):

the square root of seven is \[\sqrt{7}\] not \[\sqrt{-7}\], so 7 is not the solution.

OpenStudy (anonymous):

-7 is.

OpenStudy (anonymous):

Does anyone see this?